gr3ve3:

if your car is going at an acceleration of 5 m/s squared and it uses 5000n whats its mass?

toga:

To find the mass of the car, we can use the formula F = ma, where F is the force applied on the car, m is the mass of the car, and a is the acceleration of the car. Given that the acceleration of the car is 5 m/s^2 and the force applied is 5000 N, we can rearrange the formula to solve for m: m = F / a Substituting the given values, we get: m = 5000 N / 5 m/s^2 m = Therefore, the mass of the car is what

toga:

if you need more help tag me

gr3ve3:

@toga wrote:
To find the mass of the car, we can use the Given that the acceleration of the car is 5 m/s^2 and the force applied is 5000 N, we can rearrange the formula to solve for m: m = F / a Substituting the given values, we get: m = 5000 N / 5 m/s^2 m = Therefore, the mass of the car is what
your a life saver lil bro